OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 A. x = 9 and x = 11 B. x = −9 and x = 15 C. x = 9 and x = −15 D. No solution PLEASE help im stuck on this question!

pooja195 (pooja195):

add 4 to both sides |2x+6|-4+4=24+4 |2x+6|=24 <-----equation we are left with 2x+6=24 or 2x+6=-24 <---set solution to a postive value and a negative value since we are looking for the absolute value Solve both for x

pooja195 (pooja195):

2x+6=24 and 2x+6=-24

OpenStudy (anonymous):

thank you so much! @pooja195

pooja195 (pooja195):

:)

pooja195 (pooja195):

What did you get as the answer?

OpenStudy (anonymous):

C

pooja195 (pooja195):

Good job :-)